Use the confidence interval to find the margin of error and the sample mean. ( 1.70, 2.00 )
Answer #1

Solution :

\bar x  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = \bar x = (1.70+2.00) / 2=1.85

Margin of error = E = Upper confidence interval - \bar x   = 2.00-1.85=0.15

Margin of error = E =0.15